# Deviation Inequalities for Rényi Divergence Estimators via Variational Expression

### v1

@@ 5.1. Proof of Lemma 2
$Z_{n,\mathcal{F}_{n,\sigma}}:=\sup_{f\in\mathcal{F}_{n,\sigma}}\left|\sum_{i=1}^{n}\big{(}f(X_{i})-\mathbb{E}_{\mu}[f]\big{)}\right|\quad\mbox{ and }\quad\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}:=\sup_{f\in\mathcal{E}_{n,\sigma}}\left|\sum_{i=1}^{n}\big{(}f(X_{i})-\mathbb{E}_{\nu}[f]\big{)}\right|,$
and $\mathcal{F}_{n,\sigma}$ and $\mathcal{E}_{n,\sigma}$ are as defined in Lemma 2. By an application of Theorem 5, we obtain
$\mathbb{E}[Z_{n,\mathcal{F}_{n,\sigma}}]$ $\leq c\sqrt{n}J(1,F_{n}*\varphi_{\sigma},\mathcal{F}_{n,\sigma},\mu)\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu},$ (65a)
$\mathbb{E}[\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}]$ $\leq c\sqrt{n}J(1,e^{F_{n}}*\varphi_{\sigma},\mathcal{E}_{n,\sigma},\nu)\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{2,\nu}.$ (65b)
@@
$\mathbb{P}\left(\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\big{|}\geq 2n^{-\frac{1}{2}}t_{n,\sigma,\mu,\nu}+z\right)$
$\leq\mathbb{P}\left(\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\big{|}\geq 2n^{-1}\mathbb{E}\left[Z_{n,\mathcal{F}_{n,\sigma}}\right]+2n^{-1}\mathbb{E}\left[\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}\right]+z\right)$
$\leq\mathbb{P}\left(Z_{n,\mathcal{F}_{n,\sigma}}+\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}\geq 2\mathbb{E}\left[Z_{n,\mathcal{F}_{n,\sigma}}\right]+2\mathbb{E}\left[\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}\right]+nz\right)$
$\leq\mathbb{P}\left(Z_{n,\mathcal{F}_{n,\sigma}}\geq 2\mathbb{E}\left[Z_{n,\mathcal{F}_{n,\sigma}}\right]+\frac{nz}{2}\right)+\mathbb{P}\left(\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}\geq 2\mathbb{E}\left[\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}\right]+\frac{nz}{2}\right).$ (66)
Applying (63) to each term above and scaling $z$ by appropriate constants (depending on $q$), we obtain (24).
To show (25f)-(25g) and thus complete the proof, we will use the following maximal inequality upper bounding the Orlicz norm of maximum of finitely many random variables.
Lemma 6 (Orlicz norm of maxima).
For any $0<q<\infty$ and $\psi_{q}(x)=e^{x^{q}}-1$, we have
The proof of (67) is given in Appendix C.2, which is similar in spirit to the proof of [75, Lemma 2.2.2.], albeit without invoking convexity of $\psi_{q}$ (required here since $\psi_{q}=e^{x^{q}}-1$ is not convex near origin for $0<q<1$).
Continuing with the proof of Lemma 2, it follows from (67) that
$\left\|M_{n,\mathcal{F}_{n,\sigma}}\right\|_{\psi_{q}}$ $\leq\big{(}5\log(2n+1)\big{)}^{\frac{1}{q}}\left\|F_{n}*\varphi_{\sigma}(X)\right\|_{\psi_{q}},$
$\left\|\tilde{M}_{n,\mathcal{E}_{n,\sigma}}\right\|_{\psi_{q}}$ $\leq\big{(}5\log(2n+1)\big{)}^{\frac{1}{q}}\left\|e^{F_{n}}*\varphi_{\sigma}(Y)\right\|_{\psi_{q}},$
where RHS of the equations above are finite by the assumption $\left\|F_{n}*\varphi_{\sigma}(X)\right\|_{\psi_{q}}\vee\left\|e^{F_{n}}*\varphi_{\sigma}(Y)\right\|_{\psi_{q}}$ $\infty$ for some $q\in(0,1]$. Further, by using (103) followed by (67), we obtain
[…]
$\check{M}_{n,\mathcal{E}_{n,\sigma}}\leq\left\lceil\frac{1}{q}\right\rceil!\left\|\tilde{M}_{n,\mathcal{E}_{n,\sigma}}\right\|_{\psi_{q}}\leq\left\lceil\frac{1}{q}\right\rceil!\big{(}5\log(2n+1)\big{)}^{\frac{1}{q}}\left\|e^{F_{n}}*\varphi_{\sigma}(Y)\right\|_{\psi_{q}}.$
The proof of (26) is analogous up to (66), and the conclusion then follows by applying (62c) instead and scaling $z$ appropriately.
@@ 5.2. Proof of Lemma 3
$\qquad\qquad\qquad\qquad\qquad+\varphi_{\sigma}(x-y)\left|B_{k,\sigma}(x-y)\right|\left\|x-z\right\|\left(\frac{\left\|x-z\right\|}{2\sigma^{2}}+\frac{\left\|x-y\right\|}{\sigma^{2}}\right).$
@@
$\max_{\left|s\right|\leq\left|k\right|}\left|\tilde{c}_{k,s}\right|\leq\max_{\left|s\right|\leq\left|k\right|}\frac{|k|!}{|s|!(|k|-2|s|)!2^{|s|}}\leq\frac{|k|!}{\left(\left\lfloor\frac{\left|k\right|}{3}\right\rfloor!\right)^{2}2^{\left\lfloor\frac{\left|k\right|}{3}\right\rfloor}}\leq 3\left|k\right|^{\frac{3}{2}}(2\left|k\right|)^{\frac{\left|k\right|}{3}}.$
@@
$=\sum_{l=0}^{i-1}(u_{i-l}-v_{i-l})\left(\prod_{j=1}^{i-l-1}u_{j}\right)\left(\prod_{j=i-l+1}^{i}v_{j}\right)$
@@
$\hat{c}_{d,\left|k\right|}:=\frac{|k|!(\left|k\right|+1)^{d}}{\left(\left\lfloor\frac{\left|k\right|}{3}\right\rfloor!\right)^{2}2^{\left\lfloor\frac{\left|k\right|}{3}\right\rfloor}}.$
@@
$\qquad\qquad\qquad\left(1+\frac{\left\|x-z\right\|^{2}}{2\sigma^{2}}+\frac{\left\|x-z\right\|\left\|x-y\right\|}{\sigma^{2}}\right)+\left|B_{k,\sigma}(x-y)\right|\left(\frac{\left\|x-z\right\|}{2\sigma^{2}}+\frac{\left\|x-y\right\|}{\sigma^{2}}\right)\Bigg{)},$
@@
$\left|f_{\sigma}(x)\right|$ $\leq\int_{\mathbb{R}^{d}}\left|f(y)\right|\varphi_{\sigma}(x-y)dy$
$\leq\int_{\mathbb{R}^{d}}0.5b(1+\left\|y\right\|^{q})\varphi_{\sigma}(x-y)dy$
@@
$\leq 2^{q-1}\frac{\underaccent{\bar}{\beta}!(\underaccent{\bar}{\beta}+1)^{d}}{\left(\left\lfloor\frac{\underaccent{\bar}{\beta}}{3}\right\rfloor!\right)^{2}2^{\left\lfloor\frac{\underaccent{\bar}{\beta}}{3}\right\rfloor}}\left(\sqrt{(q+1)\underaccent{\bar}{\beta}!}\vee\sqrt{(q+\underaccent{\bar}{\beta})^{q+\underaccent{\bar}{\beta}}}\right)\left(d\vee\sqrt{d^{\underaccent{\bar}{\beta}+q+1}}\right).$ (79)
@@
Let $\mathcal{F}_{\sigma}(b,q,\mathcal{X})$ denote the class of functions obtained by restricting the domain of functions in $\mathcal{F}_{\sigma}(b,q)$ to $\mathcal{X}$. It follows from (32) that $\mathcal{F}_{\sigma}\big{(}b,q,\mathbb{B}_{d}(r)\big{)}\subseteq C_{L}^{\beta}\big{(}\mathbb{B}_{d}(r)\big{)}$ for any $\beta>0$ with $L:=\check{c}_{d,\beta,q}b(1+r^{q})\sigma^{-(\underaccent{\bar}{\beta}\vee 1)}$, where $\check{c}_{d,\beta,q}$ is given in (79). Then, applying (80) with $L$ as above and noting that $\left\|F_{b,q}*\varphi_{\sigma}\right\|_{\infty,\mathbb{B}_{d}(r)}\geq 0.5b$ yields
[…]
@@
$\qquad\qquad\qquad\left(2^{q}\frac{\underaccent{\bar}{\beta}!(\underaccent{\bar}{\beta}+1)^{d}}{\left(\left\lfloor\frac{\underaccent{\bar}{\beta}}{3}\right\rfloor!\right)^{2}2^{\left\lfloor\frac{\underaccent{\bar}{\beta}}{3}\right\rfloor}}\left(\sqrt{(q+1)\underaccent{\bar}{\beta}!}\vee\sqrt{(q+\underaccent{\bar}{\beta})^{q+\underaccent{\bar}{\beta}}}\right)\left(d\vee\sqrt{d^{\underaccent{\bar}{\beta}+q+1}}\right)\right)^{\frac{d}{\beta}}.$ (82)
@@
This proves (33a). Similarly, (81) yields (33b) by using that $\left\|F_{b,q}*\varphi_{\sigma}\right\|_{p,\gamma}\geq 0.5b$ for $p\geq 1$. Finally, the proof of the penultimate claim for $\mathcal{E}_{\sigma,\alpha}(b,q)$ follows straightforwardly via the same proof by replacing $v_{b,q}(\mathcal{X})$ by $e^{\alpha v_{b,q}(\mathcal{X})}$.
@@ 5.3. Proof of Lemma 1
where $\tilde{c}_{q}$ is the maximum of the constant appearing in [41, Theorem 6.21] and $1$. We will use the following statistical version of Talagrand’s inequality given in [6].
@@
Equation (84) follows from [6, Theorem 2.1] by setting $\eta=\frac{1+\alpha}{1-\alpha}$.
@@
By symmetrization inequality (see e.g. [75, Lemma 2.3.1]), we have
@@
$\mathbb{E}\left[\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right|\right]$ $\leq 8\mathbb{E}\left[\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right|\right]+8t_{0}(T_{n})$
$\leq 8\mathbb{E}\left[M_{n}\right]+8t_{0}(T_{n}),$
@@
From [41, Theorem 6.21], we obtain
$\left\|Z_{n}^{\prime\prime}\right\|_{\psi_{q}}\leq\tilde{c}_{q}\left(\left\|Z_{n}^{\prime\prime}\right\|_{1}+\left\|\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})-\mathbb{E}_{\mu}\left[f\mathbbm{1}_{\{F_{n}>T_{n}\}}\right]\right|\right\|_{\psi_{q}}\right).$ (92)
@@
$\left\|\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})-\mathbb{E}_{\mu}\big{[}f\mathbbm{1}_{\{F_{n}>T_{n}\}}\big{]}\right|\right\|_{\psi_{q}}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}2^{\frac{1}{q}-1}\left\|\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right|\right\|_{\psi_{q}}+2^{\frac{1}{q}-1}\left\|\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|\mathbb{E}_{\mu}\big{[}f\mathbbm{1}_{\{F_{n}>T_{n}\}}\big{]}\right|\right\|_{\psi_{q}}$
@@
$Z^{{}^{\prime}(\varepsilon)}_{n}$ $=\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i=1}^{n}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})\right|\right]$
$=\mathbb{E}_{\varepsilon_{\mathcal{I}(X^{n})}}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i\in\mathcal{I}(X^{n})}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})\right|\right]$
$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{\varepsilon_{\mathcal{I}(X^{n})}}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i\in\mathcal{I}(X^{n})}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})+\sum_{i\in\bar{\mathcal{I}}(X^{n})}\mathbb{E}_{\varepsilon_{i}}\left[\varepsilon_{i}f(X_{i})\right]\right|\right]$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i\in\mathcal{I}(X^{n})}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})+\sum_{i\in\bar{\mathcal{I}}(X^{n})}\varepsilon_{i}f(X_{i})\right|\right]$
$=\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}}\left|\sum_{i\in\mathcal{I}(X^{n})}\varepsilon_{i}f(X_{i})+\sum_{i\in\bar{\mathcal{I}}(X^{n})}\varepsilon_{i}f(X_{i})\right|\right]$

### v3

@@ 5.1. Proof of Lemma 2
$Z_{n,\mathcal{F}_{n,\phi}}:=\sup_{f\in\mathcal{F}_{n,\phi}}\left|\sum_{i=1}^{n}\big{(}f(X_{i})-\mathbb{E}_{\mu}[f]\big{)}\right|\quad\mbox{ and }\quad\tilde{Z}_{n,\mathcal{E}_{n,\phi}}:=\sup_{f\in\mathcal{E}_{n,\phi}}\left|\sum_{i=1}^{n}\big{(}f(X_{i})-\mathbb{E}_{\nu}[f]\big{)}\right|,$
and $\mathcal{F}_{n,\phi}$ and $\mathcal{E}_{n,\phi}$ are as defined in Lemma 2. By an application of Theorem 5, we obtain
$\mathbb{E}[Z_{n,\mathcal{F}_{n,\phi}}]$ $\leq c\sqrt{n}J(1,F_{n}*\phi,\mathcal{F}_{n,\phi},\mu)\left\|F_{n}*\phi\right\|_{2,\mu},$
$\mbox{and}\quad\mathbb{E}[\tilde{Z}_{n,\mathcal{E}_{n,\phi}}]$ $\leq c\sqrt{n}J(1,e^{F_{n}}*\phi,\mathcal{E}_{n,\phi},\nu)\left\|e^{F_{n}}*\phi\right\|_{2,\nu}.$
@@
$\mathbb{P}\left(\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\Phi\middle\|\hat{\nu}_{n}*\Phi\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\Phi\middle\|\nu*\Phi\right)\big{|}\geq 2n^{-\frac{1}{2}}t_{n,\varphi_{\sigma},\mu,\nu}+z\right)$
$\leq\mathbb{P}\left(\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\Phi\middle\|\hat{\nu}_{n}*\Phi\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\Phi\middle\|\nu*\Phi\right)\big{|}\geq 2n^{-1}\mathbb{E}\left[Z_{n,\mathcal{F}_{n,\phi}}\right]+2n^{-1}\mathbb{E}\left[\tilde{Z}_{n,\mathcal{E}_{n,\phi}}\right]+z\right)$
$\leq\mathbb{P}\left(Z_{n,\mathcal{F}_{n,\phi}}+\tilde{Z}_{n,\mathcal{E}_{n,\phi}}\geq 2\mathbb{E}\left[Z_{n,\mathcal{F}_{n,\phi}}\right]+2\mathbb{E}\left[\tilde{Z}_{n,\mathcal{E}_{n,\phi}}\right]+nz\right)$
$\leq\mathbb{P}\left(Z_{n,\mathcal{F}_{n,\phi}}\geq 2\mathbb{E}\left[Z_{n,\mathcal{F}_{n,\phi}}\right]+\frac{nz}{2}\right)+\mathbb{P}\left(\tilde{Z}_{n,\mathcal{E}_{n,\phi}}\geq 2\mathbb{E}\left[\tilde{Z}_{n,\mathcal{E}_{n,\phi}}\right]+\frac{nz}{2}\right).$ (61)
Applying (59) in Theorem 7 to each term above and scaling $z$ by appropriate constants (depending on $p$), we obtain (25).
The proof of (27) is analogous up to (61), and the conclusion then follows by applying (58c) instead, noting that $\sup_{f\in\mathcal{F}_{n,\phi}}\mathbb{E}_{\mu}\big{[}\big{(}f-\mathbb{E}_{\mu}[f]\big{)}^{2}\big{]}\leq\left\|F_{n}*\phi\right\|_{2,\mu}^{2}$ and scaling $z$ appropriately.
@@ 5.2. Proof of Lemma 3
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\varphi_{\sigma}(x-y)\left|B_{k,\sigma}(x-y)\right|\left\|x-z\right\|\left(\frac{\left\|x-z\right\|}{2\sigma^{2}}+\frac{\left\|x-y\right\|}{\sigma^{2}}\right).$
@@
$\max_{\left|s\right|\leq\left|k\right|}\left|\tilde{c}_{k,s}\right|\leq\max_{\left|s\right|\leq\left|k\right|}\frac{|k|!}{|s|!(|k|-2|s|)!2^{|s|}}\leq|k|!\left(\left\lfloor\frac{\left|k\right|}{3}\right\rfloor!\right)^{-2}2^{-\left\lfloor\frac{\left|k\right|}{3}\right\rfloor}\leq 3\left|k\right|^{\frac{3}{2}}(2\left|k\right|)^{\frac{\left|k\right|}{3}}.$
@@
$=\sum_{l=0}^{i-1}(u_{i-l}-v_{i-l})\left(\prod_{j=1}^{i-l-1}u_{j}\right)\left(\prod_{j=i-l+1}^{i}v_{j}\right).$
@@
$\hat{c}_{d,\left|k\right|}:=|k|!(\left|k\right|+1)^{d}\left(\left\lfloor\frac{\left|k\right|}{3}\right\rfloor!\right)^{-2}2^{-\left\lfloor\frac{\left|k\right|}{3}\right\rfloor}.$
@@
$\qquad\qquad\qquad\left(1+\frac{\left\|x-z\right\|^{2}}{2\sigma^{2}}+\frac{\left\|x-z\right\|\left\|x-y\right\|}{\sigma^{2}}\right)+\left|B_{k,\sigma}(x-y)\right|\left(\frac{\left\|x-z\right\|}{2\sigma^{2}}+\frac{\left\|x-y\right\|}{\sigma^{2}}\right)\Bigg{)}$
@@
$\left|f_{\sigma}(x)\right|\leq\int_{\mathbb{R}^{d}}\left|f(y)\right|\varphi_{\sigma}(x-y)dy$ $\leq\int_{\mathbb{R}^{d}}0.5b(1+\left\|y\right\|^{q})\varphi_{\sigma}(x-y)dy$
@@
$\leq 2^{q-1}\underaccent{\bar}{\beta}!(\underaccent{\bar}{\beta}+1)^{d}\left(\left\lfloor\frac{\underaccent{\bar}{\beta}}{3}\right\rfloor!\right)^{-2}2^{-\left\lfloor\frac{\underaccent{\bar}{\beta}}{3}\right\rfloor}\left(\sqrt{(q+1)\underaccent{\bar}{\beta}!}\vee\sqrt{(q+\underaccent{\bar}{\beta})^{q+\underaccent{\bar}{\beta}}}\right)\left(d\vee\sqrt{d^{\underaccent{\bar}{\beta}+q+1}}\right).$ (73)
@@
Let $\mathcal{F}_{\varphi_{\sigma}}(b,q,\mathcal{X})$ denote the class of functions obtained by restricting the domain of functions in $\mathcal{F}_{\varphi_{\sigma}}(b,q)$ to $\mathcal{X}$. It follows from (31) that $\mathcal{F}_{\varphi_{\sigma}}\big{(}b,q,\mathbb{B}_{d}(r)\big{)}\subseteq C_{L}^{\beta}\big{(}\mathbb{B}_{d}(r)\big{)}$ for any $\beta>0$ with $L:=\check{c}_{d,\beta,q}b(1+r^{q})\sigma^{-(\underaccent{\bar}{\beta}\vee 1)}$, where $\check{c}_{d,\beta,q}$ is given in (73). Then, applying (74) with $L$ as above and noting that $\left\|F_{b,q}*\varphi_{\sigma}\right\|_{\infty,\mathbb{B}_{d}(r)}\geq 0.5b$ yields
$\log N\big{(}\left\|F_{b,q}*\varphi_{\sigma}\right\|_{\infty,\mathbb{B}_{d}(r)}\epsilon,\mathcal{F}_{\varphi_{\sigma}}\big{(}b,q\big{)},\left\|\cdot\right\|_{\infty,\mathbb{B}_{d}(r)}\big{)}$ $\leq c_{d,\beta}(1+r)^{d}L^{\frac{d}{\beta}}\left(\left\|F_{b,q}*\varphi_{\sigma}\right\|_{\infty,\mathbb{B}_{d}(r)}\epsilon\right)^{-\frac{d}{\beta}}$
@@
$\quad\left(2^{q}\underaccent{\bar}{\beta}!(\underaccent{\bar}{\beta}+1)^{d}\left(\left\lfloor\frac{\underaccent{\bar}{\beta}}{3}\right\rfloor!\right)^{-2}2^{-\left\lfloor\frac{\underaccent{\bar}{\beta}}{3}\right\rfloor}\left(\sqrt{(q+1)\underaccent{\bar}{\beta}!}\vee\sqrt{(q+\underaccent{\bar}{\beta})^{q+\underaccent{\bar}{\beta}}}\right)\left(d\vee\sqrt{d^{\underaccent{\bar}{\beta}+q+1}}\right)\right)^{\frac{d}{\beta}}.$ (76)
@@
This proves (32a). Similarly, (75) yields (32b) by using that $\left\|F_{b,q}*\varphi_{\sigma}\right\|_{p,\gamma}\geq 0.5b$ for $p\geq 1$. Finally, the proof of the penultimate claim for $\mathcal{E}_{\varphi_{\sigma},\alpha}(b,q)$ follows straightforwardly via the same proof by replacing $v_{b,q}(\mathcal{X})$ by $e^{\alpha v_{b,q}(\mathcal{X})}$.
@@ 5.3. Proof of Lemma 1
where $\tilde{c}_{q}$ is the maximum of the constant appearing in [50, Theorem 6.21] and $1$. We will use the following statistical version of Talagrand’s inequality given in [7].
@@
Equation (78) follows from [7, Theorem 2.1] by setting $\eta=\frac{1+\alpha}{1-\alpha}$.
@@
By symmetrization inequality (see e.g. [83, Lemma 2.3.1]), we have
@@
$\mathbb{E}\left[\sup_{f\in\mathcal{F}_{n}}\left|\sum_{i=1}^{n}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right|\right]$ $\leq 8\mathbb{E}\left[\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right|\right]+8t_{0}(T_{n})\leq 8\mathbb{E}\left[M_{n}\right]+8t_{0}(T_{n}),$
@@
From [50, Theorem 6.21], we obtain
$\left\|Z_{n}^{\prime\prime}\right\|_{\psi_{q}}\leq\tilde{c}_{q}\left(\left\|Z_{n}^{\prime\prime}\right\|_{1}+\bigg{\|}\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})-\mathbb{E}_{\mu}\left[f\mathbbm{1}_{\{F_{n}>T_{n}\}}\right]\right|\bigg{\|}_{\psi_{q}}\right).$ (86)
@@
$\bigg{\|}\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})-\mathbb{E}_{\mu}\big{[}f\mathbbm{1}_{\{F_{n}>T_{n}\}}\big{]}\right|\bigg{\|}_{\psi_{q}}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}2^{\frac{1}{q}-1}\bigg{\|}\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|f\mathbbm{1}_{\{F_{n}>T_{n}\}}(X_{i})\right|\bigg{\|}_{\psi_{q}}+2^{\frac{1}{q}-1}\bigg{\|}\max_{1\leq i\leq n}\sup_{f\in\mathcal{F}_{n}}\left|\mathbb{E}_{\mu}\big{[}f\mathbbm{1}_{\{F_{n}>T_{n}\}}\big{]}\right|\bigg{\|}_{\psi_{q}}$
@@
To obtain upper bounds on $\overline{M_{n}}$ and $\left\|M_{n}\right\|_{\psi_{q}}$, we will use the following maximal inequality upper bounding the Orlicz norm of maximum of finitely many random variables.
Lemma 6 (Orlicz norm of maxima).
For any $0<q<\infty$ and $\psi_{q}(x)=e^{x^{q}}-1$, we have
The proof of (87) is given in Appendix C.2, which is similar in spirit to the proof of [83, Lemma 2.2.2.], albeit without invoking convexity of $\psi_{q}$ (required here since $\psi_{q}=e^{x^{q}}-1$ is not convex near origin for $0<q<1$).
From (87), it follows that $\left\|M_{n}\right\|_{\psi_{q}}\leq\big{(}5\log(2n+1)\big{)}^{1/q}\left\|F_{n}\right\|_{\psi_{q}}<\infty.$ Further, by using (98) given in Appendix C.1 followed by (87),
$\overline{M}_{n}\leq\left\lceil\frac{1}{q}\right\rceil!\left\|M_{n}\right\|_{\psi_{q}}\leq\left\lceil\frac{1}{q}\right\rceil!\big{(}5\log(2n+1)\big{)}^{\frac{1}{q}}\left\|F_{n}\right\|_{\psi_{q}}<\infty.$
@@
$Z^{{}^{\prime}(\varepsilon)}_{n}$ $=\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\sum_{i=1}^{n}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})\Bigg{|}\right]$
$=\mathbb{E}_{\varepsilon_{\mathcal{I}(X^{n})}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\sum_{i\in\mathcal{I}(X^{n})}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})\Bigg{|}\right]$
$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}_{\varepsilon_{\mathcal{I}(X^{n})}}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\sum_{i\in\mathcal{I}(X^{n})}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})+\sum_{i\in\bar{\mathcal{I}}(X^{n})}\mathbb{E}_{\varepsilon_{i}}\left[\varepsilon_{i}f(X_{i})\right]\Bigg{|}\right]$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\sum_{i\in\mathcal{I}(X^{n})}\varepsilon_{i}f\mathbbm{1}_{\{F_{n}\leq T_{n}\}}(X_{i})+\sum_{i\in\bar{\mathcal{I}}(X^{n})}\varepsilon_{i}f(X_{i})\Bigg{|}\right]$
$=\mathbb{E}_{\varepsilon}\left[\sup_{f\in\mathcal{F}}\Bigg{|}\sum_{i\in\mathcal{I}(X^{n})}\varepsilon_{i}f(X_{i})+\sum_{i\in\bar{\mathcal{I}}(X^{n})}\varepsilon_{i}f(X_{i})\Bigg{|}\right]$